# Contextualized Counterspeech: Strategies for Adaptation, Personalization, and Evaluation

### v2

@@ 7. Discussion and Conclusions
The difficulty at generating well-contextualized messages may be due to the need of additional information, as recent studies suggest that LLMs might struggle to combine multiple instructions and information, which can degrade their output (Giorgi et al., 2024; Beck et al., 2024). The manual analysis of some generated counterspeech, reported in Appendix C, supports this hypothesis. However, this limitation is likely to be mitigated by the adoption of larger LLMs (Hackenburg et al., 2024). Therefore, future availability of ever-larger LLMs is likely to render adapted and personalized counterspeech, and more broadly moderation interventions (Cresci et al., 2022), increasingly advantageous.
@@
Evaluation. Our results also bear important implications for the evaluation of AI-generated counterspeech. Most existing works rely on a small set of algorithmic indicators to assess the quality of the generated counterspeech (Doğanç and Markov, 2023; Hong et al., 2024; Tekiroglu et al., 2022). However, our results, alongside other recent studies (Zubiaga et al., 2024), reveal that these indicators correlate poorly with human assessment. This discrepancy suggests that indicators and human evaluators focus on different aspects of the counterspeech. Consequently, future research should adopt nuanced evaluation methods that incorporate both algorithmic and human assessments, to avoid possibly misleading conclusions.
@@
In summary, our findings indicate that human evaluators found some instances of contextualized AI-generated counterspeech particularly persuasive, highlighting the potential for AI-generated solutions to effectively address human misbehavior on online platforms. In addition, the results of our extensive evaluations call for a combined approach that integrates both algorithmic and human assessments in counterspeech evaluations. Therefore, our work points toward an increased human-AI collaboration (Pedreschi et al., 2024) that leverages the strengths of both to enhance the effectiveness of content moderation.
@@
Limitations and Future Work. Despite our extensive experimentation and evaluation, our results rely upon certain technological and methodological choices. For example, we experimented with a single LLM and we evaluated a limited set of strategies for generating adapted and personalized counterspeech. Different LLMs or alternative adaptation and personalization strategies might yield different results. Similar considerations apply to the results of our algorithmic and human evaluations. The former is limited by the quantitative indicators that we considered. These, although in line with the state-of-the-art, can gauge but a small set of characteristics. The latter is limited by the representativeness and reliability of crowdsourced evaluations. These limitations highlight the need for more research and experimentation on the effectiveness of contextualized counterspeech. Other than these, future work should explore more sophisticated models and techniques for personalizing and adapting counterspeech and should ensure that AI-generated counterspeech is unbiased and fair. Future research should investigate methods to detect possible biases in counterspeech generation and assess the long-term effects of these interventions on varied user groups.
@@ 8. Acknowledgments
This work is partially supported by the European Union – Next Generation EU within the PRIN 2022 framework project PIANO (Personalized Interventions Against Online Toxicity); by the PNRR-M4C2 (PE00000013) “FAIR-Future Artificial Intelligence Research” - Spoke 1 ”Human-centered AI”, funded under Next Generation EU; and by the Italian Ministry of Education and Research (MUR) in the framework of the FoReLab projects (Departments of Excellence).
Appendix A Prompts
@@
Appendix B Crowdsourcing questionnaire
B.1. Task description
Each participant in our crowdsourcing experiment was allowed to complete the questionnaire only once and received \$0.70 as compensation, which, given the average completion time, is above the US minimum wage. Upon providing their informed consent to take part in the experiment, participants received the following description of the task: "Your task is to evaluate a set of counterspeech responses to toxic messages posted on social media based on several criteria. With counterspeech we mean a response that addresses or challenges harmful, offensive, or toxic content with the aim to encourage a more respectful and constructive communication. Please, carefully consider the toxic post and corresponding response below, then rate the following statements from strongly disagree (1) to strongly agree (5).
B.2. Counterspeech questions
@@
  • Artificiality: The response was generated by artificial intelligence.
@@
B.3. Socio-demographic questions
@@
  • Which of the following describes best your race/ethnicity? [Asian/Asian American, Black/African American,
@@
Appendix C Counterspeech examples

### v3

@@ 7. Discussion and Conclusions
The difficulty at generating well-contextualized messages may be due to the need of additional information, as recent studies suggest that LLMs might struggle to combine multiple instructions and information, which can degrade their output (Giorgi et al., 2025; Beck et al., 2024). The manual analysis of some generated counterspeech, reported in Appendix E, supports this hypothesis. However, this limitation is likely to be mitigated by the adoption of larger LLMs (Hackenburg et al., 2024). Therefore, future availability of ever-larger LLMs is likely to render adapted and personalized counterspeech, and more broadly moderation interventions, increasingly advantageous (Cresci et al., 2022).
@@
Evaluation. Our results also bear important implications for the evaluation of AI-generated counterspeech. Most existing works rely on a small set of algorithmic indicators to assess the quality of the generated counterspeech. However, our results, alongside other recent studies (Zubiaga et al., 2024; Hengle et al., 2025), reveal that these indicators correlate poorly with human assessment. This discrepancy suggests that indicators and human evaluators focus on different aspects of the counterspeech. Consequently, future research should adopt nuanced evaluation methods that incorporate both algorithmic and human assessments, to avoid possibly misleading conclusions.
@@
In summary, our findings indicate that human evaluators found some instances of contextualized AI-generated counterspeech particularly persuasive, showing the potential for AI-generated solutions to effectively address human misbehavior on online platforms. Further, the results of our extensive evaluations call for a combined approach to integrate both algorithmic and human assessments in counterspeech evaluations. Thus, our work points toward an increased human-AI collaboration (Pedreschi et al., 2024) that leverages the strengths of both to enhance the effectiveness of content moderation.
@@
Limitations and Future Work. Despite our extensive experimentation and evaluation, our results rely upon certain technological and methodological choices. We experimented with a single LLM and we evaluated a limited set of strategies for generating adapted and personalized counterspeech. Different LLMs or alternative adaptation and personalization strategies might yield different results. Similar considerations apply to the results of our algorithmic and human evaluations. The former is limited by the quantitative indicators that we considered. These, although in line with the state-of-the-art, can gauge but a small set of characteristics. The latter is limited by the representativeness and reliability of crowdsourced evaluations. These limitations highlight the need for more research and experimentation on the effectiveness of contextualized counterspeech. Other than these, future work should explore more sophisticated models and techniques for personalizing and adapting counterspeech and should ensure that AI-generated counterspeech is unbiased and fair. Future research should investigate methods to detect possible biases in counterspeech generation and assess the long-term effects of these interventions on varied user groups.
@@ 8. Acknowledgments
This work is partially supported by the European Union – NextGenerationEU within the ERC project DEDUCE (Data-driven and User-centered Content Moderation) under grant #101113826; the PRIN 2022 project PIANO (Personalized Interventions Against Online Toxicity) under CUP B53D23013290006; and the the PNRR MUR project FAIR: Future AI Research (PE00000013). Partial support was also received by the MUR in the framework of the FoReLab projects (Departments of Excellence) and by the project “Advancing Italian Language Processing with Small-Scale Training and Preference Modeling” (IsCb8_AILP), funded by CINECA under the ISCRA initiative, for the availability of HPC resources and support.
Appendix A Summary of relevant Works
Table 2 reports the main works in the areas of LLM-generated counterspeech, LLM persuasiveness, and LLM adaptation and personalization.
[…]
Figure 7. Human evaluation results (non-contextual condition) based on answers from those participants who reported using social media “very often”. Statistical significance: ***: $p<0.01$, **: $p<0.05$.
Appendix B Counterspeech properties
This section motivates and describes all counterspeech properties introduced in Section 3.
Desired properties of effective counterspeech:
Relevant properties of contextualized AI-generated counterspeech:
Appendix C Prompts
@@
config. toxic message generated counterspeech issue BaPrHi Probably true considering the Christian god is a genocidal lunatic that condones slavery and expects misogyny. Hi there! I understand that you may have strong feelings about religion, but please remember that all beliefs and opinions deserve respect and kindness. Name-calling and spreading hate only serves to divide us and perpetuate harmful stereotypes. Instead, I invite you to share your thoughts and engage in constructive dialogue. Let’s work together to create a welcoming and inclusive community for everyone. BaPr Ahh ya sorry! Forgot about that idiot haha. Hi there! I understand that you may have forgotten about Saskatchewan, but please refrain from using derogatory language to refer to its residents. Everyone deserves respect and kindness, regardless of where they are from. Let’s focus on having a productive and inclusive conversation. Is there anything specific you would like to discuss? HsHi You don’t have to shit in everyone else’s cereal because they aren’t digesting the news exactly the way you want. I think it’s important to remember that people can have different opinions and still get along. MuRe result of this administration. Now is not the time to engage in forgive and forget healing bullshit. I’m sorry but I don’t see how your statement is hateful. inadequate MuRePrHi For a while, red tide was the slogan to counter Blue wave. Completely ignoring the fact that the comparison is to an abundance of dead fish. Can’t make this shit up. I think it’s more about the color than anything else. inadequate MuRePrHi Obama and Osama are different by two letters, B and S. BS = bullshit! They’ve been mocking us all along, in plain sight. I think you’re confusing “BS” with “BS”. meaningless MuRe Elon reinstated Trump but he hasn’t used it. Or as I call it, throwing more dog shit into the camp fire. I don’t think Elon has ever said he would reinstate Trump. incorrect MuRePrHi Won’t someone please think of the children that we accidentally forgot to murder I thought about them, but then I realized that if we didn’t kill them, they would grow up to be adults who will vote for democrats. toxic
Table 3. Examples of toxic messages and the corresponding counterspeech generated by some of the evaluated configurations. The reported examples highlight both cases of effective and problematic counterspeech.
Appendix D Crowdsourcing questionnaire
D.1. Task description
Each participant in our crowdsourcing experiment was allowed to complete the questionnaire only once and received \$0.70 as compensation, which, given the average completion time, is above the US minimum wage. Upon providing their informed consent to take part in the experiment, participants received the following description of the task: "Your task is to evaluate a set of counterspeech responses to toxic messages posted on social media based on several criteria. With counterspeech we mean a response that addresses or challenges harmful, offensive, or toxic content with the aim to encourage a more respectful and constructive communication. Consider the toxic post and corresponding response below, then rate the following statements from strongly disagree (1) to strongly agree (5).
D.2. Counterspeech questions
@@
  • Artificiality: The response was generated by AI.
@@
D.3. Socio-demographic questions
@@
  • Which of the following describes your race/ethnicity? [Asian/Asian American, Black/African American,
@@
Appendix E Counterspeech examples